OpenStudy (babtaooche):

Can someone check my answer, I'll fan and medal Determine algebraically whether the function is even, odd, or neither even nor odd. f as a function of x is equal to x plus quantity 12 over x

OpenStudy (babtaooche):

I said odd

OpenStudy (priyar):

So what is your answer?

OpenStudy (babtaooche):

I messed up in that equation!

OpenStudy (priyar):

i think you have made a typo.. in your first post

OpenStudy (babtaooche):

The equation is \[f(x)=x+\frac{ 12 }{ x }\]

OpenStudy (priyar):

Do you know what is the condition for a function to be odd?

OpenStudy (babtaooche):

It depends on the exponents, right? That's what my teacher told me @priyar

OpenStudy (priyar):

Actually there is a small condition that you can check to decide about this. If f(-x)=f(x) then even funct. If f(-x)= -f(x) then odd funct.

OpenStudy (priyar):

Now can you tell me ?

OpenStudy (babtaooche):

Is it an even function?

OpenStudy (babtaooche):

@priyar

OpenStudy (priyar):

Did you check the condition? Can your show your steps?

OpenStudy (babtaooche):

I'm confused..

OpenStudy (priyar):

It is actually simple. Let me show you an example.. if i am given f(x) = 3x i will check for f(-x), here all i need to do is replace "x" with "-x" so what will be f(-x)?

OpenStudy (babtaooche):

It would make it odd

OpenStudy (priyar):

Good!! now do the same for our Q

OpenStudy (priyar):

@babtaooche ?

OpenStudy (babtaooche):

Would it be odd?

OpenStudy (priyar):

Good job!

OpenStudy (babtaooche):

Thank you!

OpenStudy (priyar):

Now you know how to solve such Q's right?

OpenStudy (babtaooche):

Yes!

OpenStudy (priyar):

Great! you deserve a medal